Cane Company manufactures two products called Alpha and Beta that sell for $120 and $80, respectively. Each product uses only one type of raw material that costs $6 per pound. The company has the capacity to annually produce 100,000 units of each product. Its average cost per unit for each product at this level of activity are given below:

Alpha Beta
Direct materials $ 30 $ 12
Direct labor 20 15
Variable manufacturing overhead 7 5
Traceable fixed manufacturing overhead 16 18
Variable selling expenses 12 8
Common fixed expenses 15 10
Total cost per unit $ 100 $ 68

The company considers its traceable fixed manufacturing overhead to be avoidable, whereas its common fixed expenses are unavoidable and have been allocated to products based on sales dollars.

4. Assume that Cane expects to produce and sell 90,000 Betas during the current year. One of Cane’s sales representatives has found a new customer who is willing to buy 5,000 additional Betas for a price of $39 per unit. What is the financial advantage (disadvantage) of accepting the new customer's order?

5. Assume that Cane expects to produce and sell 95,000 Alphas during the current year. One of Cane’s sales representatives has found a new customer who is willing to buy 10,000 additional Alphas for a price of $80 per unit; however pursuing this opportunity will decrease Alpha sales to regular customers by 5,000 units.

a. What is the financial advantage (disadvantage) of accepting the new customer’s order?

b. Based on your calculations above should the special order be accepted?

6. Assume that Cane normally produces and sells 90,000 Betas per year. What is the financial advantage (disadvantage) of discontinuing the Beta product line?
Answer #1

Solution 4:

Contribution margin per unit on special order of Beta= Selling price - Variable cost per unit

= $39 - ($12 + $15 + $5 + $8) = -$1

If Cane accepts the customer’s offer for 5000 Beta, then increase (decrease) in operating income = 5000*(-$1) = -$5,000

Solution 5:

Regular contribution margin per unit for Alpha = $120 - ($30 + $20 + $7 + $12) = $51 per unit

Contribution margin per unit on special order of alpha= Selling price - Variable cost per unit

= $80 - ($30 + $20 + $7 + $12) = $11 per unit

If Cane accepts the customer’s offer for 10000 alpha, additional contribution margin from special order = 10000*$11 = $110,000

If Cane accept special order then loss of contribution margin on regular order = 5000*$51 = $255,000

Incremental net operating income if the order is accepted = $110,000 - $255,000 = ($145,000)

Solution 6:

Differential Analysis - Sale Beta (90000 units) (alt 1) or Discontinue Beta (Alt2)
Particulars Sale Beta (90000 Units)
(Alt 1)
Discontinue Beta (Alt 2) Differential effect on income (Alt 2)
Details Amount Details Amount
Revenue 90000*$80 $7,200,000.00 $0.00 -$7,200,000.00
Costs:
Direct Material 90000*$12 $1,080,000.00 $0.00 -$1,080,000.00
Direct Labor 90000*$15 $1,350,000.00 $0.00 -$1,350,000.00
Variable manufacturing Overhead 90000*$5 $450,000.00 $0.00 -$450,000.00
Variable Selling Expenses 90000*$8 $720,000.00 $0.00 -$720,000.00
Traceable Fixed manufacturing overhead 100000*$18 $1,800,000.00 $0.00 -$1,800,000.00
Common fixed expenses 100000*$10 $1,000,000.00 100000*$10 $1,000,000.00 $0.00
Income / (Loss) $800,000.00 -$1,000,000.00 -$1,800,000.00